# Oriented Object Detection in Optical Remote Sensing Images using Deep Learning: A Survey

### v4

@@ 8 State-of-the-Art Methods
As a comprehensive survey on oriented object detection, this paper introduces recent advances and provides a structural taxonomy based on detection frameworks, OBB representations, and Other strategies in Sections 3, 4, and 5, respectively. In this section, we select several publicly available detectors to compare them in a unified manner. Specifically, we take DOTA-V1.0 dataset since it contains almost all the typical challenges of this task, including arbitrary orientations, large-scale variations, and large aspect ratio. We report the performance of the state-of-the-art detectors in terms of mAP in Table 4. The top three scores in each category are marked in red, blue, and bold, respectively. According to the performance comparison and previous discussion, we concentrate on the key elements that evolved in oriented object detection, including detection frameworks, OBB regression, feature representations, and data augmentation.
@@
(1) Detection frameworks. Two-stage detectors achieve the best performance in terms of mAP since they can extract accurate region-based features more suited for classification and regression tasks. The typical two-stage oriented object detection methods commonly designed a rotated proposal generation scheme to obtain more accurate proposals, such as RoI Transformer (Ding et al., 2019) and Oriented RCNN (Xie et al., 2021). Similarly, the majority of one-stage detectors introduce a refined stage to align features, including RDet (Yang et al., 2021b), SANet (Han et al., 2022), CFA (Guo et al., 2021), and Oriented RepPoints (Li et al., 2022b). Benefiting from the additional refined stage and the advanced loss functions, one-stage detectors can also reach approximate accuracy to two-stage detectors. (2) OBB regression. Advanced loss functions are conducive to alleviating the problems caused by orientation parameters and achieving better regression in one-stage detectors, including GWD (Yang et al., 2021c), KLD (Yang et al., 2021d), and KFIoU (Yang et al., 2022). As can be seen in the results of multi-scale training and testing, when taking ResNet-152 as the backbone RDet-GWD, RDet-KLD, and RDet-KFIoU outperform RDet (Yang et al., 2021b) by , , and in terms of mAP, respectively. In particular, for objects with a large aspect ratio, such as bridge (BR) and harbor (HB), as well as objects with a large scale, such as ground track field (GTF) and Roundabout (RA), RDet with advanced loss functions can achieve approximate mAP improvement. The reason may be that Gaussian distribution based methods can jointly optimize the OBB parameters. Nevertheless, there is a huge gap between different OBB representations, e.g., the midpoint offset representation enable oriented RCNN to outperform all other OBB representation methods.
@@
(3) Feature representations. As one of the most important components in oriented object detection, backbone networks play a critical role in learning high-level semantic feature representation. The most widely used backbone networks include ResNet (He et al., 2016; Xie et al., 2017) series and Transformer architectures (Dosovitskiy et al., 2021; Xu et al., 2021b; Zhang et al., 2023b; Liu et al., 2021b). Although Transformer-based methods have dominated the majority of computer vision tasks, they slightly outperform CNN-based counterparts in oriented object detection. Specifically, Oriented RCNN-RVSA (Wang et al., 2022) and Oriented RCNN-STD (Yu et al., 2023), as the top-performing detectors based on Transformer, only surpasses Oriented RCNN (Xie et al., 2021) by and in terms of mAP, respetively. In addition, compared to CNNs, Transformers suffer from longer training convergence time and expensive computing costs.
(4) Data augmentation.
The multiscale training and testing (MS) generally first resize the original images to three scales (i.e., ), which are then cropped to patches of with a stride of 524. In contrast, the single scale training and testing (SS) only crops the original images to patches of with a stride of 824. As seen in Table 4, detectors with MS achieve an average approximate improvement in terms of mAP. For objects with a large scale, such as ground track field (GTF) and soccer ball field (SBF), and objects with small scale and weak features, such as helicopter (HC), MS promotes detectors by in terms of mAP. Extensive experiments have proven that MS is a useful strategy to alleviate scale variations. However, MS suffers from extremely long times for training and inference, which are about 10 times that of SS.
Method | Backbone | PL1 | BD | BR | GTF | SV | LV | SH | TC | BC | ST | SBF | RA | HA | SP | HC | mAP | |
Single-scale | ||||||||||||||||||
One-stage | RDet (Yang et al., 2021b) | R101 | 88.76 | 83.09 | 50.91 | 67.27 | 76.23 | 80.39 | 86.72 | 90.78 | 84.68 | 83.24 | 61.98 | 61.35 | 66.91 | 70.63 | 53.94 | 73.79 |
SANet (Han et al., 2022) | R101 | 88.70 | 81.41 | 54.28 | 69.75 | 78.04 | 80.54 | 88.04 | 90.69 | 84.75 | 86.22 | 65.03 | 65.81 | 76.16 | 73.37 | 58.86 | 76.11 | |
CFA (Guo et al., 2021) | R-152 | 89.08 | 83.20 | 54.37 | 66.87 | 81.23 | 80.96 | 87.17 | 90.21 | 84.32 | 86.09 | 52.34 | 69.94 | 75.52 | 80.76 | 67.96 | 76.67 | |
RDet-KLD (Yang et al., 2021d) | R50 | 88.90 | 84.17 | 55.80 | 69.35 | 78.72 | 84.08 | 87.00 | 89.75 | 84.32 | 85.73 | 64.74 | 61.80 | 76.62 | 78.49 | 70.89 | 77.36 | |
RDet-GWD (Yang et al., 2021c) | R-152 | 88.74 | 82.63 | 54.88 | 70.11 | 78.87 | 84.59 | 87.37 | 89.81 | 84.79 | 86.47 | 66.58 | 64.11 | 75.31 | 78.43 | 70.87 | 77.57 | |
Oriented RepPoints (Li et al., 2022b) | Swin-T | 89.11 | 82.32 | 56.71 | 74.95 | 80.70 | 83.73 | 87.67 | 90.81 | 87.11 | 85.85 | 63.60 | 68.60 | 75.95 | 73.54 | 63.76 | 77.63 | |
SANet-GCL (Ming et al., 2024) | Swin-T | 89.02 | 84.72 | 57.21 | 74.72 | 80.68 | 84.11 | 88.68 | 90.87 | 87.36 | 87.20 | 65.60 | 70.08 | 76.91 | 78.55 | 65.35 | 78.73 | |
Two-stage | ROI Trans. (Ding et al., 2019) | R101 | 88.64 | 78.52 | 43.44 | 75.92 | 68.81 | 73.68 | 83.59 | 90.74 | 77.27 | 81.46 | 58.39 | 53.54 | 62.83 | 58.93 | 47.67 | 69.56 |
SCRDet (Yang et al., 2019a) | R101 | 89.98 | 80.65 | 52.09 | 68.36 | 68.36 | 60.32 | 72.41 | 90.85 | 87.94 | 86.86 | 65.02 | 66.68 | 66.25 | 68.24 | 65.21 | 72.61 | |
Gliding Vertex (Xu et al., 2021a) | R50 | 89.64 | 85.00 | 52.26 | 77.34 | 73.01 | 73.14 | 86.82 | 90.74 | 79.02 | 86.81 | 59.55 | 70.91 | 72.94 | 70.86 | 57.32 | 75.02 | |
AOPG (Cheng et al., 2022a) | R101 | 89.14 | 82.74 | 51.87 | 69.28 | 77.65 | 82.42 | 88.08 | 90.89 | 86.26 | 85.13 | 60.60 | 66.30 | 74.05 | 67.76 | 58.77 | 75.39 | |
DODet (Cheng et al., 2022b) | R101 | 89.61 | 83.10 | 51.43 | 72.02 | 79.16 | 81.99 | 87.71 | 90.89 | 86.53 | 84.56 | 62.21 | 65.38 | 71.98 | 70.79 | 61.93 | 75.89 | |
SCRDet++ (Yang et al., 2022a) | R101 | 89.77 | 83.90 | 56.30 | 73.98 | 72.60 | 75.63 | 82.82 | 90.76 | 87.89 | 86.14 | 65.24 | 63.17 | 76.05 | 68.06 | 70.24 | 76.20 | |
ReDet (Han et al., 2021a) | ReR50 | 88.79 | 82.64 | 53.97 | 74.00 | 78.13 | 84.06 | 88.04 | 90.89 | 87.78 | 85.75 | 61.76 | 60.39 | 75.96 | 68.07 | 63.59 | 76.25 | |
QPDet (Yao et al., 2023) | R50 | 89.55 | 83.66 | 54.06 | 73.93 | 78.93 | 83.08 | 88.29 | 90.89 | 86.60 | 84.80 | 62.03 | 65.55 | 74.16 | 70.09 | 58.16 | 76.25 | |
Oriented RCNN (Xie et al., 2021) | R101 | 88.86 | 83.48 | 55.27 | 76.92 | 74.27 | 82.10 | 87.52 | 90.90 | 85.56 | 85.33 | 65.51 | 66.82 | 74.36 | 70.15 | 57.28 | 76.28 | |
RPGAOD (Qiao et al., 2023) | R50 | 89.34 | 83.53 | 54.39 | 76.59 | 78.09 | 81.44 | 87.64 | 90.83 | 85.89 | 85.33 | 65.44 | 64.96 | 73.73 | 70.31 | 59.50 | 76.47 | |
Oriented RCNN-ARC (Pu et al., 2023) | ARC-R101 | 89.39 | 83.58 | 57.51 | 75.94 | 78.75 | 83.58 | 88.08 | 90.90 | 85.93 | 85.38 | 64.03 | 68.65 | 75.59 | 72.03 | 65.68 | 77.70 | |
AO2-DETR (Dai et al., 2022) | R50 | 89.27 | 84.97 | 56.67 | 74.89 | 78.87 | 82.73 | 87.35 | 90.50 | 84.68 | 85.41 | 61.97 | 69.96 | 74.68 | 72.39 | 71.62 | 77.73 | |
Oriented RCN-RVSA(Wang et al., 2022) | ViTAE | 89.38 | 84.26 | 59.39 | 73.19 | 79.99 | 85.36 | 88.08 | 90.87 | 88.50 | 86.53 | 58.93 | 72.24 | 77.31 | 79.59 | 71.24 | 78.99 | |
Multi-scale | ||||||||||||||||||
One-stage | RDet (Yang et al., 2021b) | R-152 | 89.80 | 83.77 | 48.11 | 66.77 | 78.76 | 83.27 | 87.84 | 90.82 | 85.38 | 85.51 | 65.67 | 62.68 | 67.53 | 78.56 | 72.62 | 76.47 |
RDet-DCL (Yang et al., 2021a) | R-152 | 89.26 | 83.60 | 53.54 | 72.76 | 79.04 | 82.56 | 87.31 | 90.67 | 86.59 | 86.98 | 67.49 | 66.88 | 73.29 | 70.56 | 69.99 | 77.37 | |
SANet (Han et al., 2022) | R50 | 88.89 | 83.60 | 57.74 | 81.95 | 79.94 | 83.19 | 89.11 | 90.78 | 84.87 | 87.81 | 70.30 | 68.25 | 78.30 | 77.01 | 69.58 | 79.42 | |
RDet-GWD (Yang et al., 2021c) | R-152 | 89.66 | 84.99 | 59.26 | 82.19 | 78.97 | 84.83 | 87.70 | 90.21 | 86.54 | 86.85 | 73.47 | 67.77 | 76.92 | 79.22 | 74.92 | 80.23 | |
RDet-KLD (Yang et al., 2021d) | R-152 | 89.92 | 85.13 | 59.19 | 81.33 | 78.82 | 84.38 | 87.50 | 89.80 | 87.33 | 87.00 | 72.57 | 71.35 | 77.12 | 79.34 | 78.68 | 80.63 | |
RDet-KFIoU (Yang et al., 2022) | R-152 | 88.89 | 85.14 | 60.05 | 81.13 | 81.78 | 85.71 | 88.27 | 90.87 | 87.12 | 87.91 | 69.77 | 73.70 | 79.25 | 81.31 | 74.56 | 81.03 | |
SANet-GCL (Ming et al., 2024) | Swin-T | 89.40 | 85.19 | 60.39 | 79.88 | 81.37 | 85.80 | 88.99 | 90.85 | 85.73 | 88.02 | 70.38 | 73.52 | 83.56 | 82.49 | 75.63 | 81.41 | |
Two-stage | SCRDet++ (Yang et al., 2022a) | R101 | 90.05 | 84.39 | 55.44 | 73.99 | 77.54 | 71.11 | 86.05 | 90.67 | 87.32 | 87.08 | 69.62 | 68.90 | 73.74 | 71.29 | 65.08 | 76.81 |
AO2-DETR (Dai et al., 2022) | R50 | 89.95 | 84.52 | 56.90 | 74.83 | 80.86 | 83.47 | 88.47 | 90.87 | 86.12 | 88.55 | 63.21 | 65.09 | 79.09 | 82.88 | 73.46 | 79.22 | |
ReDet (Han et al., 2021a) | ReR50 | 88.81 | 82.48 | 60.83 | 80.82 | 78.34 | 86.06 | 88.31 | 90.87 | 88.77 | 87.03 | 68.65 | 66.90 | 79.26 | 79.71 | 74.67 | 80.10 | |
ReDet-DEA (Liang et al., 2022) | ReR50 | 89.92 | 83.84 | 59.65 | 79.88 | 80.11 | 87.96 | 88.17 | 90.31 | 88.93 | 88.46 | 68.93 | 65.94 | 78.04 | 79.69 | 75.78 | 80.37 | |
DODet (Cheng et al., 2022b) | R50 | 89.96 | 85.52 | 58.01 | 81.22 | 78.71 | 85.46 | 88.59 | 90.89 | 87.12 | 87.80 | 70.50 | 71.54 | 82.06 | 77.43 | 74.47 | 80.62 | |
AOPG (Cheng et al., 2022a) | R50 | 89.88 | 85.57 | 60.90 | 81.51 | 78.70 | 85.29 | 88.85 | 90.89 | 87.60 | 87.65 | 71.66 | 68.69 | 82.31 | 77.32 | 73.10 | 80.66 | |
Oriented RCNN (Xie et al., 2021) | R50 | 89.84 | 85.43 | 61.09 | 79.82 | 79.71 | 85.35 | 88.82 | 90.88 | 86.68 | 87.73 | 72.21 | 70.80 | 82.42 | 78.18 | 74.11 | 80.87 | |
RoI Trans.-KFIoU (Yang et al., 2022) | Swin-T | 89.44 | 84.41 | 62.22 | 82.51 | 80.10 | 86.07 | 88.68 | 90.90 | 87.32 | 88.38 | 72.80 | 71.95 | 78.96 | 74.95 | 75.27 | 80.93 | |
QPDet (Yao et al., 2023) | R50 | 90.14 | 85.31 | 60.98 | 79.92 | 80.21 | 85.04 | 88.80 | 90.87 | 86.45 | 88.04 | 70.88 | 71.72 | 82.99 | 80.55 | 73.19 | 81.00 | |
RPGAOD (Qiao et al., 2023) | R50 | 89.93 | 85.00 | 60.16 | 81.58 | 79.63 | 84.94 | 88.71 | 90.90 | 86.24 | 87.65 | 72.94 | 72.61 | 81.24 | 79.78 | 76.76 | 81.20 | |
Oriented RCNN-RVSA(Wang et al., 2022) | ViTAE | 88.97 | 85.76 | 61.46 | 81.27 | 79.98 | 85.31 | 88.30 | 90.84 | 85.06 | 87.50 | 66.77 | 73.11 | 84.75 | 81.88 | 77.58 | 81.24 | |
Oriented RCNN-STD (Yu et al., 2023) | HiViT-B | 89.15 | 85.03 | 60.79 | 82.06 | 80.90 | 85.76 | 88.45 | 90.83 | 87.71 | 87.29 | 73.99 | 71.25 | 85.18 | 82.17 | 82.95 | 82.24 |
1
PL: plane. BD: baseball diamond. BR: bridge. GTF: ground track field. SV: small vehicle. LV: large vehicle. SH: ship. TC: tennis court. BC: baseball court. ST: storage tank. SBF: soccer ball field. RA: roundabout.
HA: harbor. SP: swimming pool. HC: helicopter. -
2
The results of these methods are the best-reported results from corresponding papers. indicates anchor-free methods.

### v5

@@ 8 State-of-the-Art Methods
As a comprehensive survey on oriented object detection, this paper introduces recent advances and provides a structural taxonomy based on detection frameworks, OBB representations, and other strategies in Sections 3, 4, and 5, respectively. In this section, we select several publicly available detectors to compare them in a unified manner. Specifically, we take DOTA-V1.0 dataset since it contains almost all the typical challenges of this task, including arbitrary orientations, large-scale variations, and large aspect ratio. We report the performance of the state-of-the-art detectors in terms of mAP and show the crucial modules of each detector in Tab. 5. According to the performance comparison and previous discussion, we concentrate on the key elements that evolved in oriented object detection, including detection frameworks, OBB regression, feature representations, and common issues.
@@
(1) Detection Frameworks. Two-stage detectors achieve the best performance in terms of mAP since they can extract accurate region-based features more suited for classification and regression tasks. The typical two-stage oriented object detection methods commonly designed a rotated proposal generation scheme to obtain more accurate rotated proposals, such as RoI Transformer [29] and Oriented RCNN [174]. Similarly, the majority of one-stage and anchor-free detectors introduce a refined stage to align features, including R3Det [193], S2ANet [47], CFA [38, 40], and Oriented RepPoints [74]. Benefiting from the additional refined stage and the advanced loss functions, one-stage detectors can also reach approximate accuracy to two-stage detectors. Despite achieving State-of-the-Art performance in general object detection, DETR-based methods still lag behind other competitors in oriented object detection, even with more training epochs. This may be because the query paradigm could not cover rotated objects adequately. To this end, it is desirable to further investigate DETR-based methods to compete in this field.
(2) OBB Regression. Advanced loss functions are conducive to alleviating the problems caused by orientation parameters and achieving better regression, including Gaussian distance based loss (e.g., GWD [198], KLD [200], KFIoU [203]). These methods draw upon the trigonometric encoder and joint optimization to achieve strong performance. On the other hand, there is a huge gap between different OBB representations, e.g., the midpoint offset representation enable Oriented RCNN to outperform Rotated Faster RCNN by approximately 2 mAP in single-scale and multi-scale results. In addition, novel OBB representation methods elegantly avoid angular boundary discontinuity thus enhancing model performance, but they rely on complex post-processing and additional modules, including customized loss functions (e.g., CIoU [38]) or assigner (e.g., APAA [74]). Therefore, advanced loss functions and OBB representations are crucial for improving regression accuracy.
@@
(3) Feature Representations. As one of the most important components in oriented object detection, backbone networks play a critical role in learning high-level semantic feature representation. The most widely used backbone networks include ResNet [65, 177] series and Transformer architectures [22, 181, 224, 85]. Although Transformer-based methods dominate the field of computer vision tasks, and significantly outperform CNN-based counterparts in oriented object detection, achieving state-of-the-art performance. Specifically, Oriented RCNN-RVSA [169] and Oriented RCNN-STD [191], as the top-performing detectors based on Transformer, surpass Oriented RCNN [174] by and in terms of mAP, respectively. Nevertheless, compared to CNNs, Transformers suffer from longer training convergence time and expensive computing costs.
(4) Common Issues. Attention mechanism and semantic mask modules are effective ways to reduce background noise and enhance object information, e.g., SCRDet [199], TSH [197], HRDet [182]. Regarding the issue of large aspect ratios, the re-weighted loss function and novel assignment strategies can effectively improve detection accuracy, e.g., RPGAOD [123], DFDet [173], CFL [144], CGCDet [170]. These components will be removed during the inference process, thus they will not affect the inference speed. On the other hand, as seen in Tab. 5, detectors with MS444The multi-scale training and testing (MS) generally first resize the original images to three scales (i.e., ), which are then cropped to patches of with a stride of 524. In contrast, the single-scale training and testing (SS) only crop the original images to patches of with a stride of 824. achieve an average approximate improvement in terms of mAP, proving that MS is a useful strategy to alleviate scale variations. However, MS suffers from extremely long times for training and inference, which are about 10 times that of SS. All in all, addressing common issues like background noise, large aspect ratios, and scale variations is crucial for improving oriented object detection.
Method | Publication | Baseline | Backbone | RPN | Assigner | Head | Reg Loss | Cls Loss | OBB Representation | FE | mAP-SS | mAP-MS |
Two-stage | ||||||||||||
Rotated Faster RCNN [130] | NIPS 2015 | Faster RCNN | Res50 | – | – | – | – | – | – | 73.40∗ | 78.49∗ | |
RoI Transformer [29] | CVPR 2019 | Faster RCNN | Res50 | RRoI Learner | – | – | – | – | – | 75.63∗ | 80.43∗ | |
Oriented RCNN [174] | ICCV 2021 | Faster RCNN | Res50 | Oriented RPN | – | – | – | – | Midpoint Offset | 75.69∗ | 80.02∗ | |
Gliding Vertex [176] | TPAMI 2021 | Faster RCNN | Res50 | – | – | – | – | – | Gliding Vertex | 75.02 | ||
ReDet [48] | CVPR 2021 | RoI Transformer | ReRes50 | – | – | – | – | – | – | 76.25 | 80.10 | |
OSKDet [86] | CVPR 2022 | Faster RCNN | Res50 | – | – | Anchor Free | – | GFL v2 | Unordered Keypoints | 76.37 | 80.91 | |
AOPG [19] | TGRS 2022 | Faster RCNN | Res50 | Rotated FCOS | – | – | – | – | – | 75.22 | 80.66 | |
DEA [80] | TGRS 2022 | ReDet | ReRes50 | FCOS+RPN | – | – | – | – | – | 80.37 | ||
RVSA [169] | TGRS 2022 | Oriented RCNN | ViT+RVSA | – | – | – | – | – | – | 78.61 | 80.80 | |
RPGAOD [123] | TGRS 2023 | Oriented RCNN | Res50 | GRG-RPN | – | – | AAO | – | – | 76.47 | 81.20 | |
QPDet [183] | TGRS 2023 | Oriented RCNN | Res50 | – | – | – | – | – | Quadrant Point | 76.25 | 81.00 | |
FRIoU Loss [124] | TGRS 2023 | Oriented RCNN | Res50 | – | – | – | FRIoU | – | – | 76.45 | 80.78 | |
CGCDet [170] | TNNLS 2024 | RoI Transformer | Res50 | – | OCP-Guided | – | CGC | – | – | 77.34 | 80.70 | |
STD [191] | AAAI 2024 | Oriented RCNN | ViT | – | – | MAEB | – | – | – | 82.24 | ||
One-stage | ||||||||||||
Rotated RetinaNet [78] | ICCV 2017 | RetinaNet | Res50 | – | – | – | – | – | 68.80∗ | |||
SCRDet [199] | ICCV 2019 | RetinaNet | SF-Net | – | – | IoU-Smooth L1 | – | – | MDA | 72.61 | ||
CSL [192] | ECCV 2020 | RetinaNet | Res50 | – | – | – | – | CSL | 69.51∗ | |||
R3Det [193] | AAAI 2021 | RetinaNet | Res50 | – | FRM | – | – | – | 70.18∗ | |||
DCL [187] | CVPR 2021 | R3Det | Res50 | – | – | – | – | DCL | 71.21 | |||
DAL [115] | AAAI 2021 | RetinaNet | Res50 | DAS | – | – | – | – | 71.44 | |||
FoRDet [236] | TGRS 2021 | R2SSD | Res50 | – | – | FARL | – | FRM | 71.44 | |||
GWD [198] | ICML 2021 | R3Det | Res50 | – | – | GWD | – | Gaussian | 71.56 | |||
KLD [200] | NeurIPS 2021 | R3Det | Res50 | – | – | KLD | – | Gaussian | 71.73 | |||
BCD [202] | TPAMI 2022 | R3Det | Res50 | – | – | BCD | – | Gaussian | 72.22 | |||
S2ANet [47] | TGRS 2022 | RetinaNet | Res50 | – | FAM | – | – | – | 74.12 | 79.42 | ||
CFC-Net [112] | TGRS 2022 | RetinaNet | Res50 | DAS | RARM | – | – | – | PAM | 73.50 | ||
DCFL [175] | CVPR 2023 | S2ANet | Res50 | DCFL | – | – | – | – | 74.26 | |||
KFIoU [203] | ICLR 2023 | R3Det | Res50 | – | – | KFIoU | – | Gaussian | 71.60 | |||
TCD [223] | TGRS 2023 | RetinaNet | Res50 | TCA | TCH | – | TCL | – | 75.18 | 80.05 | ||
FADL-Net [31] | TII 2024 | RetinaNet | Res50 | GADL | – | – | JLRQ | – | 74.80 | 79.97 | ||
TIR-Net [97] | TGRS 2024 | S2ANet | Res50 | – | SRK+RFR | – | – | – | 75.23 | 80.63 | ||
CFL [144] | TIM 2024 | S2ANet | Res50 | STS | CFS | – | – | – | 75.35 | |||
Anchor-Free | ||||||||||||
DRN [120] | CVPR 2020 | CenterNet | H104 | – | DRH | – | – | – | FSM | 70.70 | ||
O2-DNet [166] | ISPRS 2021 | FCOS | H104 | – | – | – | – | Middle lines | 72.80 | |||
CFA [38, 40] | CVPR 2021 | RepPoints | Res101 | – | CFA | CIoU | – | Convex-Hull | 75.05 | |||
Oriented RepPoints [74] | CVPR 2022 | RepPoints | Res50 | APAA | – | – | – | Adaptive Points | 75.97 | |||
GFL [158] | TGRS 2022 | CenterNet | Res50 | – | – | GF-CSL | – | CSL | 74.68 | 77.54 | ||
SASM [56] | AAAI 2022 | RepPoints | Res50 | SA-S | – | SA-M | – | 74.92 | 77.19 | |||
DHRec [116] | TPAMI 2023 | FCOS | Res50 | – | – | – | – | DHRec | 74.57 | 78.97 | ||
DRDet [216] | TGRS 2023 | FCOS | Res50 | – | OFE | – | – | DRLs | 74.85 | 79.34 | ||
TSH [197] | TGRS 2023 | FCOS | Res50 | – | – | JQE | PBL | ISM | 77.18 | |||
HRDet [182] | TCSVT 2024 | FCOS | Res50 | – | – | EIoU | – | – | HMP | 74.11 | 78.90 | |
DFDet [173] | TGRS 2024 | FCOS | Res50 | – | – | PIAS | – | – | CDMN | 74.71 | 80.37 | |
DETR-Based | ||||||||||||
AO2-DETR [25] | TCSVT 2023 | Deformable DETR | Res50 | – | AOPR | – | – | – | 70.91 | |||
ARS-DETR [207] | TGRS 2024 | DINO | Res50 | ARM | RDA | ARA-CSL | – | – | 73.79 |